# Distributed Omniscient Observers for Multi-Agent Systems

### v1

@@ Abstract
This article proposes fully distributed omniscient observers for both heterogeneous and homogeneous linear multi-agent systems, through which each agent can estimate the states of all agents. The proposed observers not only contribute to distributed Nash equilibrium seeking in multi-player games, but also provide a designable self-organization mechanism for artificial swarms to emulate biological social behaviors, including sheepdog herding and honeybee dance communication.
I Introduction
@@
As consensus-reaching has been relatively well studied, recent focus has been increasingly placed on more advanced forms of cooperation in MAS. However, the aforementioned distributed setup poses challenges to achieving advanced collaboration. Specifically, if an agent only has access to limited local information, it may fail to effectively cooperate with others for a global objective of the MAS. This motivates the development of distributed omniscient observers in this article, which aim to provide each agent with sufficient global information to support autonomous decision-making.
In prior research on leader–follower MAS, distributed observers have been commonly designed either for each follower to estimate the leader’s state [11, 13, 20, 26, 27, 21, 22], or for each agent to reconstruct its own absolute state using relative output information [29]. Different from the above approaches, the distributed omniscient observers proposed in this article enable every agent to estimate the global MAS state (which includes both its own state and the states of all other agents), while allowing each agent’s input to be persistently nonzero. This equips each agent with a “global view”, enabling the MAS to cooperatively perform complex tasks that extend beyond basic consensus. The proposed observer design method mainly builds on the framework developed in [4]. However, the method in this article emphasizes exploiting relative (neighbor-to-neighbor) output information in MAS, which can be easier and cheaper to realize in practice compared with using the absolute one.
@@
II Review of the Distributed Observers Without Using Global Inputs
II-A Communication Graph
@@
II-B Revisiting the Design Method
@@
III Distributed Omniscient Observer Design
@@
III-A Design for Heterogeneous Multi-Agent Systems
@@
III-B Design for Homogeneous Multi-Agent Systems
It follows directly that the distributed omniscient observers developed in Section III-A can be applied to the following homogeneous MAS:
@@ III-B Design for Homogeneous Multi-Agent Systems
The dynamics of the overall MAS take the same form as (1a), where $x={\rm{col}}(\breve{x}_{i})_{i=1}^{N}$, $u={\rm{col}}(\breve{u}_{i})_{i=1}^{N}$, $A=I_{N}\otimes\breve{A}$, and $B=I_{N}\otimes\breve{B}$. Moreover, (12b) aligns with (1b), where $C_{i}=\breve{C}\left({{\bar{\cal L}}_{i}}\otimes{I_{\breve{n}}}\right)$. In this case, we construct an analytical solution of the following form:
${B_{i}}$ $=\frac{{T_{id}}{\breve{B}}}{\left\|{{{\bar{\cal L}}_{i}}}\right\|}$ (14a)
@@
$T_{id}^{\top}$ $=\frac{{{\bar{\cal L}}_{i}}\otimes I_{\breve{n}}}{\left\|{{\bar{\cal L}}_{i}}\right\|}$ (14c)
@@
${F_{i}}$ $=-\frac{{\breve{L}_{i}}}{\left\|{{{\bar{\cal L}}_{i}}}\right\|},\ i\in\mathcal{N},$ (14e)
where ${\bar{\cal L}}_{i}$ is defined in (13), and $\breve{L}_{i}$ is chosen such that $\breve{A}+\breve{L}_{i}\breve{C}$ has eigenvalues with negative real parts. It can be verified that the above analytical solution satisfies constraints (2c)-(2g). The following two lemmas guarantee that constraints (2a) and (2b) are also satisfied.
Lemma 2.
@@
See Sections VII-A and VII-B for the proof of Lemma 2 and Lemma 3, respectively. Given the fact that choosing $u_{i}$ as in (12a) and $u_{-i}$ as $u$ preserves (5), the following result is straightforward from Lemma 1.
@@
III-C Further Extension
To implement the observers designed in Section III-B, each agent needs neighbors’ input information. In what follows, we develop distributed omniscient observers that do not require neighbors’ inputs for homogeneous MAS (11). This is achieved at the expense of undermining adaptability to the change of communication graphs, since a unified scalar gain depending on the global graph should be shared by each agent. Specifically, let us consider the distributed observers proposed in [29]:
@@ III-C Further Extension
See Section VII-C for the proof of Theorem 3.
IV Application A: Distributed Nash Equilibrium Seeking in Multi-Player Games
@@
Algorithm (20) requires each player to have real-time access to global action profile $x$. For the sake of scalability, however, communications may only occur between neighboring agents in MAS. In the case where only neighbors’ actions are directly available, the distributed omniscient observers developed in Section III can be used to implement algirithm (20), by providing each player with an estimate of the global action profile. Specifically, the distributed Nash equilibrium seeking algorithm is designed as follows:
@@
See Section VII-D for the proof of Theorem 4. Further discussions on Theorem 4 are as follows:
@@
V Application B: Self-Organized Social Behavior Emergence in Artificial Swarms
@@
V-A Confine Companions to a Convex Hull
@@
Figure 1: Agent trajectories (start/end positions denoted by circles/crosses) in Section V-A.
V-B Summon Companions by Circling
The second example is inspired by the dancing behaviors of honeybees — they use dance language to communicate the location and the abundance of nectar sources to other members of the hive. In this example, the leader agents will circle at different places with different speeds, which simulates the behavior of the dancing honeybees. The follower agents will assemble at the leaders’ places respectively, according to the ratio of their circling speeds. Followers tend to be attracted to leaders with higher speeds.
@@ V-B Summon Companions by Circling
Figure 2: Snapshots of agent positions in Section V-B.
Figure 3: Agent positions vs. time in Section V-B.
V-C Simulation Results and Discussions
Figs. 1-3 demonstrate the simulation results of the two examples formulated in Section V-A and V-B, respectively. The agents that have access to their own positions are agents 2 and 7 in the first example, and agents 2, 8, and 11 in the second example. In addition, the three leaders’ circling speeds are designed to be $v^{*}_{3}=2$, $v^{*}_{6}=3$, and $v^{*}_{9}=4$. Other detailed settings are not listed here for the sake of brevity. The following are several interpretations and discussions of the simulation results.
@@ V-C Simulation Results and Discussions
  • Fig. 1 shows that all followers can move into the convex hull formed by the leaders. Since the leaders are allowed to move freely, the shape and position of the convex hull can be time-varying.
@@
  • Fig. 3 shows that two, three, and four followers are attracted to the three leaders respectively, in proportion to the circling speeds of the leaders. Different initial positions of the agents may lead to a different final result, but will preserve the ratio of the number of followers that each leader attracts.
  • In both examples, each follower can identify who are the leaders autonomously. This is achieved by estimating the augmented identity state (in the first example) and the velocity (in the second example) of agents.
  • Such leader-identification mechanisms can accommodate the change of leaders. Suppose in the second example that agents 3 and 6 keep circling at their speeds, while agent 9 stops circling after $t=20$s. Then, for agent 9 and its four followers, two of them will head toward position $(1,1)$, and the other three will go to $(5,5)$ autonomously.
VI Conclusion
In this article, distributed omniscient observers are proposed for heterogeneous and homogeneous linear MAS, respectively. The observer design for the latter is based on mostly relative, as well as a small amount of absolute, local input-output information of the agents. Knowledge of the global communication graph can obviate the relative input information required for the design, and vice versa. As a result, each agent can estimate the states of itself and other agents. The state estimation errors can converge to zero even though the inputs of other agents are persistently nonzero. An application in distributed Nash equilibrium seeking, and two bio-inspired simulation examples show that the proposed distributed omniscient observers can contribute to the emergence of collective intelligence in MAS.
VII Appendix
VII-A Proof of Lemma 2
@@ VII-B Proof of Lemma 3
$T_{id}^{\top}\left({{I_{N}}\otimes\breve{B}}\right)\!=\!\begin{cases}{\left\|{{\bar{\cal L}}_{i}}\right\|^{-1}}\!\begin{bmatrix}l_{i1}\breve{B}&l_{i2}\breve{B}&\cdots&l_{iN}\breve{B}\end{bmatrix},\ i\in\mathcal{N}\setminus\mathcal{R}\\[5.0pt] {\left\|{{\bar{\cal L}}_{i}}\right\|^{-1}}\!\begin{bmatrix}0_{\breve{n}\times(i-1)\breve{m}}&\breve{B}&0_{\breve{n}\times(N-i)\breve{m}}\end{bmatrix},\ i\in\mathcal{R}.\end{cases}$
Therefore, ${\rm{Im}}\breve{B}={\rm{Im}}\left[{T_{id}^{\top}\left({{I_{N}}\otimes\breve{B}}\right)}\right]$ and Lemma 3 follows.
VII-C Proof of Theorem 3
It follows from [29] that $\left\|{{{\hat{\breve{x}}}_{i}}(t)-{\breve{x}}_{i}(t)}\right\|$, i.e., $\left\|T_{id}^{\top}\left(\hat{x}_{i}(t)-x(t)\right)\right\|$, exponentially converges to zero, $\forall i\in\mathcal{N}$. Then it can be proved as in [4] that adaptive gains remain bounded and $\mathop{\lim}\nolimits_{t\to\infty}\left\|T_{iu}^{\top}\left(\hat{x}_{i}(t)-x(t)\right)\right\|=0,\ \forall i\in\mathcal{N}$.
VII-D Proof of Theorem 4
@@
$V_{o}$ $=\frac{1}{2}\varepsilon_{u}^{\top}{\left[{T_{u}^{\top}({\cal L}\otimes I){T_{u}}}\right]^{-1}}{\varepsilon_{u}}+\sum\nolimits_{i=1}^{N}{\frac{1}{{2{\phi_{i}}}}{{({\gamma_{i}}-{\gamma^{*}})}^{2}}}$
@@
$\dot{V}=$ $-\left(x-x^{*}\right)^{\top}{\rm{col}}\left[\nabla_{\breve{x}_{i}}J_{i}(x)+\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})-\nabla_{\breve{x}_{i}}J_{i}(x)\right]_{i=1}^{N}+\dot{V}_{o}$
$\leq$ $-\mu\left\|x-x^{*}\right\|^{2}\!+\left(x-x^{*}\right)^{\top}\!\!{\rm{col}}\!\left[\nabla_{\breve{x}_{i}}J_{i}(x)-\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})\right]_{i=1}^{N}\!\!+\dot{V}_{o}$
$\leq$ $-\frac{\mu}{2}\left\|x-x^{*}\right\|^{2}+\frac{1}{2\mu}\left\|{\rm{col}}\left[\nabla_{\breve{x}_{i}}J_{i}(x)-\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})\right]_{i=1}^{N}\right\|^{2}+\dot{V}_{o}$

### v2

@@ Abstract
This paper proposes distributed omniscient observers for both heterogeneous and homogeneous linear multi-agent systems, such that each agent can correctly estimate the states of all agents. The observer design is based on local input-output information available to each agent, and knowledge of the global communication graph among agents is not necessarily required. The proposed observers can contribute to distributed Nash equilibrium seeking in multi-player games and the emergence of self-organized social behaviors in artificial swarms. Simulation results demonstrate that artificial swarms can emulate animal social behaviors, including sheepdog herding and honeybee dance-based navigation.
keywords:
Multi-agent systems; consensus; distributed state estimation; fully distributed design; swarm intelligence; collective intelligent behaviors.
thanks: This research is supported by the National Research Foundation, Singapore, and PUB, Singapore’s National Water Agency under its RIE2025 Urban Solutions and Sustainability (USS) (Water) Centre of Excellence (CoE) Programme, awarded to Nanyang Environment & Water Research Institute (NEWRI), Nanyang Technological University (NTU), Singapore. This research is also supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RG95/24). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore and PUB, Singapore’s National Water Agency. The material in this paper was not presented at any conference. Corresponding author: Xunyuan Yin.
1 Introduction
@@
As consensus-reaching has been relatively well studied, recent focus has been increasingly placed on more advanced forms of cooperation in MAS. However, the aforementioned distributed setup poses challenges to achieving advanced collaboration. Specifically, if an agent only has access to limited local information, it may fail to effectively cooperate with others for a global objective of the MAS. This motivates the development of distributed omniscient observers in this paper, which aim to provide each agent with sufficient global information to support autonomous decision-making.
In prior research on leader–follower MAS, distributed observers have been commonly designed either for each follower to estimate the leader’s state [11, 13, 20, 26, 27, 21, 22], or for each agent to reconstruct its own absolute state using relative output information [29]. Different from the above approaches, the distributed omniscient observers proposed in this paper enable every agent to estimate the global MAS state (which includes both its own state and the states of all other agents), while allowing each agent’s input to be persistently nonzero. This equips each agent with a “global view”, enabling the MAS to cooperatively perform complex tasks that extend beyond basic consensus. The proposed observer design method mainly builds on the framework developed in [4]. However, the method in this paper emphasizes exploiting relative (neighbor-to-neighbor) output information in MAS, which can be easier and cheaper to realize in practice compared with using the absolute one.
The rest of the paper is organized as follows. Section 2 reviews a recent development of distributed observers for general linear time-invariant systems, where local inputs and outputs are used to estimate the global system states. Based on Section 2, Section 3 proposes distributed omniscient observers for general linear MAS. In particular, Section 3 shows that relative inputs and outputs between neighboring agents can be used to estimate the global states of homogeneous MAS. Section 4 and Section 5 demonstrate applications of the proposed distributed omniscient observers and validate effectiveness of the observers by numerical examples. Section 6 concludes this paper.
@@
2 Review of the Distributed Observers Without Using Global Inputs
2.1 Communication Graph
@@
2.2 Revisiting the Design Method
@@
3 Distributed Omniscient Observer Design
@@
3.1 Design for Heterogeneous Multi-Agent Systems
@@
3.2 Design for Homogeneous Multi-Agent Systems
It follows directly that the distributed omniscient observers developed in Section 3.1 can be applied to the following homogeneous MAS:
@@ 3.2 Design for Homogeneous Multi-Agent Systems
The dynamics of the overall MAS take the same form as (1a), where $x={\rm{col}}(\breve{x}_{i})_{i=1}^{N}$, $u={\rm{col}}(\breve{u}_{i})_{i=1}^{N}$, $A=I_{N}\otimes\breve{A}$, and $B=I_{N}\otimes\breve{B}$. Moreover, (12b) aligns with (1b), where $C_{i}=\breve{C}\left({{\bar{\mathcal{L}}}_{i}}\otimes{I_{\breve{n}}}\right)$. In this case, we construct an analytical solution of the following form:
${B_{i}}$ $=\frac{{T_{id}}{\breve{B}}}{\left\|{{{\bar{\mathcal{L}}}_{i}}}\right\|}$ (14a)
@@
$T_{id}^{\top}$ $=\frac{{{\bar{\mathcal{L}}}_{i}}\otimes I_{\breve{n}}}{\left\|{{\bar{\mathcal{L}}}_{i}}\right\|}$ (14c)
@@
${F_{i}}$ $=-\frac{{\breve{L}_{i}}}{\left\|{{{\bar{\mathcal{L}}}_{i}}}\right\|},\ i\in\mathcal{N},$ (14e)
where ${\bar{\mathcal{L}}}_{i}$ is defined in (13), and $\breve{L}_{i}$ is chosen such that $\breve{A}+\breve{L}_{i}\breve{C}$ has eigenvalues with negative real parts. It can be verified that the above analytical solution satisfies constraints (2c)-(2g). The following two lemmas guarantee that constraints (2a) and (2b) are also satisfied.
Lemma 3.1.
@@
See Sections 7.1 and 7.2 for the proof of Lemma 3.1 and Lemma 3.2, respectively. Given the fact that choosing $u_{i}$ as in (12a) and $u_{-i}$ as $u$ preserves (5), the following result is straightforward from Lemma 1.
@@
3.3 Further Extension
To implement the observers designed in Section 3.2, each agent needs neighbors’ input information. In what follows, we develop distributed omniscient observers that do not require neighbors’ inputs for homogeneous MAS (11). This is achieved at the expense of undermining adaptability to the change of communication graphs, since a unified scalar gain depending on the global graph should be shared by each agent. Specifically, let us consider the distributed observers proposed in [29]:
@@ 3.3 Further Extension
See Section 7.3 for the proof of Theorem 3.
4 Application A: Distributed Nash Equilibrium Seeking in Multi-Player Games
@@
Algorithm (20) requires each player to have real-time access to global action profile $x$. For the sake of scalability, however, communications may only occur between neighboring agents in MAS. In the case where only neighbors’ actions are directly available, the distributed omniscient observers developed in Section 3 can be used to implement algirithm (20), by providing each player with an estimate of the global action profile. Specifically, the distributed Nash equilibrium seeking algorithm is designed as follows:
@@
See Section 7.4 for the proof of Theorem 4. Further discussions on Theorem 4 are as follows:
@@
5 Application B: Self-Organized Social Behavior Emergence in Artificial Swarms
@@
5.1 Confine Companions to a Convex Hull
@@
Figure 1: Agent trajectories (start/end positions denoted by circles/crosses) in Section 5.1.
Figure 2: Agent positions vs. time in Section 5.1.
Figure 3: Adaptive gains vs. time in Section 5.1.
Figure 4: Estimation error norms vs. time in Section 5.1, where $x={\rm{col}}\left(\breve{x}_{i}\right)_{i=1}^{8}$.
5.2 Summon Companions by Circling
The second example is inspired by the dancing behaviors of honeybees — they use dance language to communicate the location and the abundance of nectar sources to other members of the hive. In this example, the leader agents will circle at different places with different speeds, which emulates the behavior of the dancing honeybees. The follower agents will assemble at the leaders’ places respectively, according to the ratio of their circling speeds. Followers tend to be attracted to leaders with higher speeds.
@@ 5.2 Summon Companions by Circling
Figure 5: Snapshots of agent positions in Section 5.2.
Figure 6: Agent positions vs. time in Section 5.2.
Figure 7: Agent velocities vs. time in Section 5.2.
Figure 8: Adaptive gains vs. time in Section 5.2.
Figure 9: Estimation error norms vs. time in Section 5.2, where $x={\rm{col}}\left(\breve{x}_{i}\right)_{i=1}^{12}$.
5.3 Simulation Results and Discussions
Figs. 1-9 demonstrate the simulation results of the two examples formulated in Section 5.1 and 5.2, respectively. The agents that have access to their own positions are agents 2 and 7 in the first example, and agents 2, 8, and 11 in the second example. In addition, the three leaders’ circling speeds are designed to be $v^{*}_{3}=2$, $v^{*}_{6}=3$, and $v^{*}_{9}=4$. Other detailed settings are not listed here for the sake of brevity. The following are several interpretations and discussions of the simulation results.
@@ 5.3 Simulation Results and Discussions
  • Figs. 1 and 2 show that all followers can move into the convex hull formed by the leaders. Since the leaders are allowed to move freely, the shape and position of the convex hull can be time-varying.
@@
  • Figs. 6 and 7 show that two, three, and four followers are attracted to the three leaders respectively, in proportion to the circling speeds of the leaders. Different initial positions of the agents may lead to a different final result, but will preserve the ratio of the number of followers that each leader attracts.
  • In both examples, each follower can identify which agents are the leaders autonomously. This is achieved by estimating the augmented identity state (in the first example) and the velocity (in the second example) of agents.
  • Fig. 2 shows that the followers keep still at the beginning. This is because the identity states of the leaders have not reached threshold $z^{*}_{t}$, and the state estimates have not reached a desired level of accuracy. Therefore, the followers cannot identify the leaders correctly in the first few seconds.
  • The leader-identification mechanisms can accommodate the change of leaders. Suppose in the second example that agents 3 and 6 keep circling at their speeds, while agent 9 stops circling after $t=20$s. Then, for agent 9 and its four followers, two of them will head toward position $(1,1)$, and the other three will go to $(5,5)$ autonomously.
  • Figs. 4 and 9 show that the state estimation errors of the distributed omniscient observers converge to zero, even though the inputs of agents are persistently nonzero. Figs. 3 and 8 show that the adaptive gains in the distributed omniscient observers remain bounded.
6 Conclusion
In this paper, distributed omniscient observers are proposed for heterogeneous and homogeneous linear MAS, respectively. The observer design for the latter is based on mostly relative, as well as a small amount of absolute, local input-output information of the agents. Knowledge of the global communication graph can obviate the relative input information required for the design, and vice versa. As a result, each agent can estimate the states of itself and other agents. The state estimation errors can converge to zero even though the inputs of other agents are persistently nonzero. An application in distributed Nash equilibrium seeking, and two bio-inspired simulation examples show that the proposed distributed omniscient observers can contribute to the emergence of collective intelligence in MAS.
7 Appendix
7.1 Proof of Lemma 3.1
@@ 7 Appendix
$T_{id}^{\top}\left({{I_{N}}\otimes\breve{B}}\right)$
$=\begin{cases}{\left\|{{\bar{\mathcal{L}}}_{i}}\right\|^{-1}}\!\begin{bmatrix}l_{i1}\breve{B}&l_{i2}\breve{B}&\cdots&l_{iN}\breve{B}\end{bmatrix},\ i\in\mathcal{N}\setminus\mathcal{R}\\[5.0pt] {\left\|{{\bar{\mathcal{L}}}_{i}}\right\|^{-1}}\!\begin{bmatrix}0_{\breve{n}\times(i-1)\breve{m}}&\breve{B}&0_{\breve{n}\times(N-i)\breve{m}}\end{bmatrix},\ i\in\mathcal{R}.\end{cases}$
Therefore, ${\rm{Im}}\breve{B}={\rm{Im}}\left[{T_{id}^{\top}\left({{I_{N}}\otimes\breve{B}}\right)}\right]$ and Lemma 3.2 follows.
7.3 Proof of Theorem 3
It follows from [29] that $\left\|{{{\hat{\breve{x}}}_{i}}(t)-{\breve{x}}_{i}(t)}\right\|$, i.e.,
$\left\|T_{id}^{\top}\left(\hat{x}_{i}(t)-x(t)\right)\right\|,$
exponentially converges to zero, $\forall i\in\mathcal{N}$. Then it can be proved as in [4] that adaptive gains remain bounded and $\mathop{\lim}\nolimits_{t\to\infty}\left\|T_{iu}^{\top}\left(\hat{x}_{i}(t)-x(t)\right)\right\|=0,\ \forall i\in\mathcal{N}$.
7.4 Proof of Theorem 4
@@
$V_{o}$ $=\frac{1}{2}\varepsilon_{u}^{\top}{\left[{T_{u}^{\top}({\mathcal{L}}\otimes I){T_{u}}}\right]^{-1}}{\varepsilon_{u}}+\sum\nolimits_{i=1}^{N}{\frac{1}{{2{\phi_{i}}}}{{({\gamma_{i}}-{\gamma^{*}})}^{2}}}$
@@
$\dot{V}=$ $-\left(x-x^{*}\right)^{\top}{\rm{col}}\left[\nabla_{\breve{x}_{i}}J_{i}(x)+\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})-\nabla_{\breve{x}_{i}}J_{i}(x)\right]_{i=1}^{N}$
$+\dot{V}_{o}$
$\leq$ $-\mu\left\|x-x^{*}\right\|^{2}\!+\left(x-x^{*}\right)^{\top}\!\!{\rm{col}}\!\left[\nabla_{\breve{x}_{i}}J_{i}(x)-\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})\right]_{i=1}^{N}$
$+\dot{V}_{o}$
$\leq$ $-\frac{\mu}{2}\left\|x-x^{*}\right\|^{2}+\frac{1}{2\mu}\left\|{\rm{col}}\left[\nabla_{\breve{x}_{i}}J_{i}(x)-\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})\right]_{i=1}^{N}\right\|^{2}$
$+\dot{V}_{o}$